What volume of air containing 21% of oxygen is required to completely burn 50g of sulfur containing 4% incombustible matter


1
Expert's answer
2021-08-05T12:43:02-0400

Mass of sulfur =10050×4​=2


Mass of combustible matter"= 50-2 = 48"


"S+O_2--->SO_2"


"O_2= \frac{22.4×48}{32}=33.6"


Volume required"=\frac {100×33.6}{21}=160L"